OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. 12 2.6 8 4

mathslover (mathslover):

\[\huge{6=3(x-2)}\] \[\huge{6=3x-6}\] \[\huge{12=3x}\] \[\huge{4=x}\]

OpenStudy (a_clan):

6 = 3(x-2) 6= 3x - 6 12= 3x 4=x

OpenStudy (anonymous):

THNKZ

OpenStudy (maheshmeghwal9):

\[\huge{\color{red}{{6=3(x-2)}}}\]\[\huge{\color{blue}{6=3x-6}}\]\[\huge{\color{blue}{{12=3x}}}\]\[\huge{\color{green}{{4=x.}}} \color{purple}{\text{Ans....}}\]Very Good @mathslover .

mathslover (mathslover):

thanks @maheshmeghwal9

OpenStudy (maheshmeghwal9):

yw:)